OpenStudy (anonymous):

Find the slope of the line that passes through the pair of points. F(–6, 8), P(–6, –5) A. 0 B. –4 C. -12/13 D. undefined Is the correct answer C?

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (anonymous):

so its C? @pooja195

TheSmartOne (thesmartone):

How did you get C??

OpenStudy (anonymous):

I did 8-(-5) and it gave me 13 @TheSmartOne

pooja195 (pooja195):

-6-6----v -6+6

OpenStudy (anonymous):

Oh so its what then?

pooja195 (pooja195):

Solve it what is -6+6?

OpenStudy (anonymous):

0

pooja195 (pooja195):

0 divided by -13 =?

OpenStudy (anonymous):

0 so the answer is A?

OpenStudy (astrophysics):

No, because that's the x - coordinates, you're dividing over 0 hence undefined.

OpenStudy (anonymous):

so its undefined?

pooja195 (pooja195):

Yes

OpenStudy (anonymous):

its A or D?

OpenStudy (astrophysics):

\[\frac{ -5-8 }{ -6-(-6) } = \frac{ -13 }{ 0 } = undefined \]

OpenStudy (anonymous):

oh okay so its D?

OpenStudy (astrophysics):

yeah..

pooja195 (pooja195):



OpenStudy (anonymous):

Okay thanks can u help me with 1 more?

OpenStudy (astrophysics):

no

TheSmartOne (thesmartone):

lolol

pooja195 (pooja195):

o.O i can help

TheSmartOne (thesmartone):

just make another post :)

OpenStudy (astrophysics):

Naw I'm just messing haha go ahead ask your question

pooja195 (pooja195):

xD

OpenStudy (anonymous):

Find the rate of change for the situation. You run 7 miles in one hour and 21 miles in three hours. A. 3 miles per hour B. 3 hours C. 7 miles D. 7 miles per hour

OpenStudy (astrophysics):

rate of change = slope

OpenStudy (astrophysics):

same formula

OpenStudy (anonymous):

Can u set up the equation for me?

OpenStudy (astrophysics):

You should try, what is rate mean, is it miles/time or time/miles what makes sense?

OpenStudy (triciaal):

the first one should be A when zero on top undefined; when zero on the bottom slope is zero

OpenStudy (anonymous):

miles/times?

OpenStudy (astrophysics):

Correct and no the first one is not 0 @triciaal

OpenStudy (anonymous):

wait what? @triciaal so its A not D?

TheSmartOne (thesmartone):

It's D :P

OpenStudy (anonymous):

So u would divide 7 by 21?

OpenStudy (astrophysics):

No, \[\frac{ y_2-y_1 }{ x_2-x_1 }\] you basically have two points

OpenStudy (anonymous):

so its D?

OpenStudy (astrophysics):

Bingo

OpenStudy (anonymous):

Thanks but the 1st one what is the correct answer then?

OpenStudy (astrophysics):

Try the math yourself

OpenStudy (anonymous):

No the first question one person said A and u said D whats the correct answer?

OpenStudy (astrophysics):

You should be confirming it

OpenStudy (astrophysics):

You're dividing over 0, which is undefined

OpenStudy (astrophysics):

Google what happens when you divide over 0

OpenStudy (astrophysics):

see if you get 0 undefined candy or a blackhole

OpenStudy (anonymous):

okay just making sure...thanks

OpenStudy (anonymous):

yeah ur right its undefined

OpenStudy (astrophysics):

:)